# Physics-informed deep learning for infectious disease forecasting

### v1

@@ 2.3 Evaluation metrics
We used two metrics to assess predictive performance: mean absolute scaled error (MASE) [56] and weighted interval score (WIS). MASE was calculated by dividing the mean absolute error (MAE) of the model prediction by the MAE of a naive model.
@@
The naive model uses the data from the past weeks as predictions for the following weeks. Specially speaking, when the forecast horizon is 1 week, the naive model uses the data of the past 1 week as the predictions for the following 1 week. When the forecast horizon is 2 weeks, the naive model uses the data of the past 2 weeks as the forecasting of the following 2 weeks. MASE smaller than 1.0 means that our PINNs model predicts more accurately than the naive model. And the lower the MASE, the more accurate the predictions.
The typical use case for infectious disease forecasting requires probabilistic predictions. For error propagation, we assume that the prediction distribution is Gaussian with mean $\mu_{t}$ equal to the point forecast for time $t$ and standard deviation $\sigma$ equal to the standard deviation of all prediction errors, constant across time yielding $\mu_{t}+\sigma\sqrt{2}{erf}^{-1}(2p-1)$ for prediction quantiles at $t$.
To evaluate probabilistic predictions with WIS, we used the ‘interval-scoring’ Python package (https://github.com/adrian-lison/interval-scoring), based on  [50], to calculate WIS, with quantile bins as required by the COVID-19 Forecast Hub, i.e., $\alpha 1=0.02$, $\alpha 2=0.05$, $\alpha 3=0.1$,…,$\alpha 11=0.9$ and quantiles including 0.010, 0.025, 0.050, 0.100, 0.150, 0.200, 0.250, 0.300, 0.350, 0.400, 0.450, 0.500, 0.550, 0.600, 0.650, 0.700, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.990.
@@ 2.4 Ablation study
Compared with traditional neural networks, we employed epidemiology information in the networks by adding a loss of ODE into the overall loss function. To investigate the impact of embedding the epidemiological domain knowledge into the neural network, we also trained a traditional feed-forward neural network (FNN) and compared its performance with our PINNs model. Both PINNs and FNNs were fitted five times with different initial configurations. Predictions were obtained as the mean of the five results.
Compared with traditional mathematical methods, we used an FNN and embedded the compartment model into the FNN. To examine the advantage of PINNs over the compartment model, we compared GISST’s performance with our PINN model. GISST[22] solves the same compartmental model by semi-parametric mathematical methods.
It is noted that early history data may not reflect the dynamics of the current type of virus due to the mutation and even negatively affect future predictions. Thus, we also examined the impact of the size of the training data set on the prediction accuracy. We first trained PINNs with all historical data. We trained PINNs with data from the past few weeks instead of all the historical data and compared their predicted performances. The results are reported in Appendix 5.1.
Finally, we also investigated applying L2 regularization to our model to avoid overfitting. L2 Regularization was achieved by adding a new term, which consists of the sum of the squared values of the network’s hyperparameters, to the original loss function of PINNs during training. The regularization term penalizes certain model parameters and adjusts them to minimize the total loss of the PINNs. We examined the impact of L2 regularization to investigate how this feature impacts the performance of the model. The results are reported in Appendix 5.2.
@@ 3.1 Point forecasts on the number of cases, deaths, and hospitalizations for the following 1 - 4 weeks
Figure 3: PINNs’ point predictions on the number of cases, deaths, and hospitalizations for the following 1 - 4 weeks.
At each time point $t$, we trained PINNs with historical data from July 2020 to $t$ and made predictions for the following 1-4 weeks. We plotted the predictions in Fig. 3, which demonstrates that PINNs predictions are consistent with the target and more accurate than the naive model for all forecasting horizons from 1 week to 4 weeks. It also shows that one-week-ahead predictions are more accurate than predictions with 2, 3, and 4-week time horizons. As the time horizon for forecasting becomes greater, the predicted values deviate more from the true values, and there are more fluctuations in the curves as well.
Evaluation metrics are reported in Table 3. In the column labeled ’Method’, the name before the dash denotes the approach, while the number is the number denotes the length of forecast horizon. Scores in red denote the model with the lowest score for a given forecast horizon in that column, e.g., PINNs - 1 week denotes PINNs predictions for the following 1 week. We find that PINNs’ MASE scores are all less than 1.0, meaning that PINNs outperform the naive model for all the forecast horizons. Among the three state variables, MASE scores of deaths are higher than the scores of cases and hospitalizations because deaths are associated with larger fluctuations in the original dataset.
@@
PINN - 2 week 0.77 0.92 0.62 29570 140 562 0.71 0.96 0.66
PINN - 3 week 0.81 0.86 0.68 43331 174 951 0.76 0.90 0.78
PINN - 4 week 0.85 0.89 0.74 57471 230 1416 0.85 0.94 0.91
@@ 3.2 Quantile forecasts on the number of cases, deaths, and hospitalizations for the following 1 - 4 weeks
Figure 4: PINNs’ quantile predictions on the number of cases, deaths, and hospitalizations for the following 1 - 4 weeks. The red dots represent the ground truth; the blue points denote the PINNs predictions; the blue region represents the uncertainty. Different polygons cover a four-week prediction interval with 1-, 2-, 3- and 4-week predictions from a time one week prior to the polygon.
@@
Similarly, at each time point $t$, we trained PINNs with historical data from July 2020 to $t$ and made predictions for the following 1-4 weeks. In this case, we plotted the continuous predictions for the following 1-4 weeks as well as the corresponding uncertainty in Fig. 4. For each group of predictions for the following 1-4 weeks, the points (denoting predictions at quantile = 0.5) gradually deviate from the target, and the light blue area (denotes uncertainty) grows larger. The longer the forecast horizons, the lower the accuracy and the larger the uncertainty.
WIS scores are reported in Table 3. Scaled WIS was computed by dividing the WIS by the WIS of naive models of the same forecast horizon. We find that PINNs WIS scores are mostly less than 1.0, which means that PINNs outperform the naive model for all the forecast horizons. The only exception is the quantile predictions by PINNs for the following 1 week.
3.3 Comparison with traditional neural network and traditional mathematical method
@@
We compared our PINNs predictions with results obtained from the traditional neural network model (FNN) without incorporating the compartment model as well as with the results of a mathematical model, named Gaussian infection state space with time dependence (GISST) [21]. As shown in Table 3, both MASE and WIS scores for PINNs are lower than the corresponding FNN scores for each forecast horizon, which demonstrates the absolute advantage of PINNs over traditional neural networks for both point forecasting and quantile forecasting. This is likely due to the regularizing effect of the compartment model. Additionally, PINNs typically outperformed GISST when forecasting hospitalizations, GISST outperformed PINNs for cases, and forecasting performance for deaths depended on whether the evaluation was for point forecasts (GISST superior) or quantile forecasts (PINNs superior). These results demonstrate the important role of epidemiological domain knowledge as a constraint to avoid unreasonable predictions that may arise in purely data-driven modeling.
@@ 4 Discussion and Summary
Epidemic forecasting can be generally considered an inverse-forward problem, meaning that we need to first infer the parameters of the model before making a forecast. In many cases, we partially know the mechanisms underlying disease transmission, which is represented by the compartment models in the form of ODEs, but some of the model parameters need to be identified by fitting them to observation data. Once the parameters are estimated, compartment models can be used to predict future disease trends and assess potential shifts induced by changes in public health policies or emerging variants. The main shortcomings of the traditional mathematical models are the limitations due to model misspecification, such as assuming transition rates are fixed, making it difficult to adapt to variants with different transmission rates. On the other hand, data-driven neural network models have emerged as a new tool for tackling complex social problems over the last several years. However, conventional neural network models require a large amount of training and labeled data to ensure good performance, and they also fail to account for domain knowledge. Furthermore, ML models may provide unrealistic predictions in cases of extrapolation as they do not take into account the mechanisms underlying relevant physical processes.
In this work, we employed PINNs to leverage the epidemiological knowledge contained in compartment models together with observed data to improve the accuracy of infectious disease forecasts. Our results have demonstrated that the proposed model possesses several features that may contribute to its superior performance. First, the PINN model utilizes NNs to capture complex and non-linear relationships within infectious disease data because neural networks can represent complex non-linear dynamics. Second, PINN models integrate domain-specific knowledge by incorporating the compartment models into the loss function of the NNs to ensure that the learned model adheres to the fundamental principles and constraints of disease dynamics. In addition, incorporating compartment models could prevent the model from overfitting, which often occurs when training machine learning models with only data.
@@
Additionally, PINNs can be used to estimate key epidemiological parameters, such as transmission rates or recovery rates, directly from the data, thereby enhancing the model’s ability to adapt to changes in the disease dynamics over time. Particularly, PINNs treat these parameters as the output of the neural network such that these parameters can be time-dependent, representing the change of the dynamics of the transmission induced by, i.e., the mutation of the virus. We also show that PINNs can continuously adapt to newly acquired data, allowing the model to continuously improve its forecasting capabilities as more information becomes available during an ongoing outbreak. PINNs can integrate data from various sources, including social, environmental, and demographic factors, providing a holistic approach to infectious disease forecasting.
@@
Recently, several Physics-Informed Neural Network (PINNs) models have been developed for infectious disease forecasting such as COVID-19 [57, 58, 59] and measles [60]. Compared with these models, our proposed model incorporates a more complex compartmental framework with nine state variables into the neural network, presenting a greater challenge for the model to align with the physics described by the ODE system. In particular, our model achieves better metrics, namely Normalized Root Mean Squared Error (NR1 and NR2) and Normal Deviation (ND) as defined in [58], than the Epidemiologically-Informed Neural Networks (EINNs) [58] on predicting the death in COVID-19, the only predicted state variables of EINNs. Additionally, while previous models primarily provide point forecasts, our approach facilitates quantile forecasts, which convey uncertainty and enhance the reliability of predictions.
Notwithstanding its capabilities of dealing with noisy data, the vanilla PINN model is not equipped with built-in mechanisms to assess the uncertainty of model predictions in response to noisy input data. One popular approach for solving regression problems with uncertainty is Gaussian processes regression (GPR). Built upon the Bayesian framework, GPR can quantify uncertainty in the solution of PDEs, but the applications of GPR have been limited to linear problems with small datasets as its computational cost increases substantially with the growing size of the dataset. One possible approach to resolve this challenge is to employ Bayesian PINNs as has been proposed by [37]. These models take advantage of the function of PINNs in solving ill-posed problems, i.e., with under-defined boundary conditions, and the capability of the Bayesian framework to quantify the uncertainty for the solution. Thus, the performance of Bayesian PINNs on infectious disease forecasting needs to be explored in a future study.
@@
This work was supported by National Institute of Health grants R21HL168507 and NSF SCH Award Number:2406212. High-performance computing resources were provided by the Center for Computation and Visualization at Brown University and The Georgia Advanced Computing Resource Center (GACRC) at the University of Georgia.

### v2

@@ 2.3 Evaluation metrics
We used two metrics to assess predictive performance: the mean absolute scaled error (MASE) [61] for point forecasting and the weighted interval score (WIS) for quantile forecasting. MASE was calculated by dividing the mean absolute error (MAE) of the model’s predictions by the MAE of a naive model of the same forecast horizon.
@@
The naive model uses data from the previous weeks as predictions for the following weeks. Specifically, when the forecast horizon is 1 week, the naive model uses data from the past week to predict the following week. Similarly, for a 2-week forecast horizon, the naive model uses data from the past 2 weeks to predict the subsequent 2 weeks. A MASE value smaller than 1.0 indicates that our PINNs model outperforms the naive model in terms of predictive accuracy. Moreover, a lower MASE value corresponds to more accurate predictions.
In infectious disease forecasting, quantile predictions, namely probabilistic predictions, are typically required to account for error propagation. We assume that the prediction distribution follows a Gaussian distribution, with mean $\mu_{t}$ equal to the point forecast at time $t$, and standard deviation $\sigma$, which represents the standard deviation of all prediction errors and is constant across time. The prediction quantiles at time $t$ are then given by $\mu_{t}+\sigma\sqrt{2}\,\text{erf}^{-1}(2p-1)$, where $\text{erf}^{-1}$ denotes the inverse error function and $p$ represents the desired quantile.
To evaluate probabilistic predictions using the WIS, we employed the ‘interval-scoring’ Python package (https://github.com/adrian-lison/interval-scoring), based on [55]. We calculated the WIS with quantile bins as required by the COVID-19 Forecast Hub, specifically $\alpha_{1}=0.02$, $\alpha_{2}=0.05$, $\alpha_{3}=0.1$, …, $\alpha_{11}=0.9$, and quantiles including 0.010, 0.025, 0.050, 0.100, 0.150, 0.200, 0.250, 0.300, 0.350, 0.400, 0.450, 0.500, 0.550, 0.600, 0.650, 0.700, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, and 0.990.
@@ 2.4 Ablation study
In contrast to traditional neural networks, we incorporated epidemiological knowledge into our model by adding the ODE loss to the overall loss function. To assess the impact of embedding domain-specific epidemiological knowledge into the neural network, we also trained a traditional neural network (denoted as NN) and compared its performance with our PINNs model. The NN shares the same architecture as the neural network component in the upper left part of Fig. 1A, but it is purely data-driven and does not integrate the compartmental model.
In contrast to traditional mathematical methods, we employed a neural network and integrated it with the compartmental model. To evaluate the advantages of PINNs over the compartmental model, we compared the performance of our PINNs with that of the Gaussian Infection State Space with Time Dependence (GISST) model [23]. GISST is a conventional mathematical model that solves the same compartmental model.
It is important to note that early historical data may not accurately capture the dynamics of current virus variants due to mutations, and may even negatively impact future predictions. To assess this, we investigated the effect of the training data window size on prediction accuracy. While PINNs were originally trained using all available historical data, we also trained models using only recent data and compared their predictive performance. The results are presented in Appendix 5.1.
Additionally, we conducted an analysis of the impact of L2 regularization on model performance and evaluated its necessity in our forecasting tasks. This analysis helps to assess whether regularization significantly contributes to improving generalization and preventing overfitting. The results of these experiments, including performance comparisons with and without L2 regularization, are presented in Appendix 5.2.
@@ 3.1 Point forecasts on the number of cases, deaths, and hospitalizations for the following 1 - 4 weeks
Figure 3: PINNs’ point predictions on the number of cases, deaths, and hospitalizations for the following 1 - 4 weeks. GISST represents a mathematical model named Gaussian infection state space with time dependence [22]. The naive model uses data from the previous weeks as predictions for the following weeks.
At each time point $t$, we trained PINNs using historical data from July 2020 up to $t$, and generated forecasts for the subsequent 1 to 4 weeks. The predictions, shown in Fig. 3, demonstrate that the outputs from PINNs are consistently aligned with the observed target values and are more accurate than those from the naive model across all forecasting horizons. Additionally, PINNs outperform most of the corresponding forecasts produced by GISST. The results also indicate that one-week-ahead predictions are the most reliable, with accuracy gradually declining for forecasts on the subsequent 2-, 3-, and 4-week horizons. As the forecasting horizon increases, the predicted values tend to diverge more from the ground truth, accompanied by greater fluctuations in the predicted curves.
Quantitative comparisons are reported in Tab. 3. In the column labeled ‘Method’, the name before the dash indicates the forecasting approach, while the number denotes the forecast horizon. Scores highlighted in red represent the best-performing model (the lowest score) for a given forecast horizon within each metric column. For example, ‘PINNs - 1 week’ refers to PINNs predictions for the subsequent one week, while ‘PINNs - 2 week’ refers to forecasts for the second week ahead. We observe that all of PINNs’ MASE scores are no more than 1.0, indicating that PINNs consistently outperform the naive model across all forecast horizons. Among the three state variables, the MASE scores for deaths are generally higher than those for cases and hospitalizations, likely due to greater variability and fluctuations in the original death data.
@@
PINN - 2 week 0.77 0.92 0.62 32954 146 556 0.79 1.00 0.66
PINN - 3 week 0.81 0.86 0.68 49272 176 882 0.87 0.91 0.72
PINN - 4 week 0.85 0.89 0.74 63311 217 1232 0.94 0.89 0.79
@@
RNN - 1 week 0.91 1.28 0.70 17653 126 278 0.76 1.27 0.64
RNN - 2 week 0.92 1.17 0.82 33544 168 642 0.81 1.15 0.76
RNN - 3 week 0.94 1.09 0.94 46641 205 1105 0.82 1.06 0.90
RNN - 4 week 0.92 0.99 1.03 53964 242 1554 0.80 0.99 1.00
LSTM - 1 week 1.00 1.29 0.79 19622 124 325 0.85 1.25 0.74
LSTM - 2 week 1.00 1.39 0.93 35249 194 786 0.85 1.33 0.93
LSTM - 3 week 0.98 1.16 1.03 49661 226 1249 0.87 1.17 1.02
LSTM - 4 week 0.94 1.09 1.04 56474 268 1624 0.84 1.09 1.05
GRU - 1 week 0.89 1.22 0.77 16952 115 324 0.73 1.16 0.74
GRU - 2 week 0.90 1.24 0.88 33573 178 749 0.81 1.22 0.88
GRU - 3 week 0.93 1.14 0.98 47267 225 1198 0.83 1.17 0.98
GRU - 4 week 0.90 1.03 1.01 55414 262 1608 0.82 1.07 1.04
Transformer - 1 week 0.96 1.14 0.81 22155 111 356 0.96 1.12 0.81
Transformer - 2 week 0.99 0.98 0.86 40494 145 730 0.97 0.99 0.86
Transformer - 3 week 1.05 0.97 0.94 56804 193 1179 1.00 1.00 0.96
Transformer - 4 week 1.08 0.91 1.04 69297 235 1629 1.03 0.96 1.05
@@ 3.2 Quantile forecasts on the number of cases, deaths, and hospitalizations for the following 1 - 4 weeks
Figure 4: PINNs’ quantile predictions on the number of cases, deaths, and hospitalizations for the following 1 - 4 weeks. The red dots represent the ground truth. The blue points indicate PINNs’ predictions and the light blue region represents the associated uncertainty in the forecasts. Each polygon spans a continuous four-week prediction interval, encompassing the 1-, 2-, 3-, and 4-week forecasts made from a time point one week prior to the start of the polygon
@@
Similarly, at each time point $t$, we trained PINNs using historical data from July 2020 up to $t$ and generated predictions for the subsequent 1 to 4 weeks. In this case, we visualized the continuous forecasts along with the associated uncertainty in Fig. 4. For each group of 1- to 4-week predictions, the points (representing the median predictions at quantile = 0.5) progressively deviate from the ground truth as the forecast horizon extends. Correspondingly, the light blue region, which represents predictive uncertainty, expands with longer horizons. This pattern highlights a common trend: as the forecast horizon increases, prediction accuracy decreases while uncertainty grows.
Quantitative comparisons of WIS are presented in Tab. 3. The scaled WIS was computed by dividing the WIS of each model by the WIS of the naive model for the corresponding forecast horizon. We observe that PINNs achieve scaled WIS scores that are generally below 1.0, indicating superior performance over the naive baseline across most forecasting horizons. The only exception occurs in the 1-week-ahead quantile predictions for death, where the PINNs’ WIS score slightly exceeds 1.0. However, this deviation is minor and still significantly lower than the scores of other models.
3.3 Comparison with sequence deep learning models and traditional mathematical model
@@
We first compared our PINNs predictions with results obtained from the traditional neural network model (NN) without incorporating the compartmental model as well as with the results of a traditional mathematical model, named GISST [22]. As shown in Tab. 3, both MASE and WIS scores for PINNs are lower than the corresponding scores of NN for each forecast horizon, which demonstrates the absolute advantage of PINNs over traditional neural networks for both point forecasting and quantile forecasting. This is likely due to the regularizing effect of the compartmental model.
Additionally, PINNs typically outperformed GISST when forecasting hospitalizations, GISST outperformed PINNs for cases, and forecasting performance for deaths depended on whether the evaluation was for point forecasts (GISST superior) or quantile forecasts (PINNs superior). These results demonstrate the important role of epidemiological domain knowledge as a constraint to avoid unreasonable predictions that may arise in purely data-driven modeling.
To further evaluate the performance of PINNs, we compared them against several state-of-the-art neural network architectures commonly used for sequential data processing. Specifically, we implemented RNNs [62], LSTMs [63], GRUs [64], and Transformer models [65] to address the same forecasting task. The RNN, LSTM, and GRU models share a similar architecture configuration. Each model consists of an input layer with three neurons, two hidden layers with 50 neurons each, and an output layer with three neurons. The models take weekly data on cases, deaths, and hospitalizations as input features and generate corresponding predictions for each of the three variables. A window size of seven weeks was used for input. Training was conducted using the Adam optimizer with a learning rate of $1\times 10^{-3}$ over 10,000 epochs.
The Transformer model is composed of an input layer with three neurons, followed by two encoder layers, each with a hidden dimensionality of 64 and four attention heads, and an output layer with three neurons. Similar to the other models, it takes weekly data on cases, deaths, and hospitalizations as input features and outputs corresponding predictions for each variable. A window size of seven weeks was used to form the input sequence. The model was trained using the Adam optimizer with a learning rate of $1\times 10^{-3}$ for 10,000 epochs.
Figure 5: A MASE comparison of Recurrent Neural Networks (RNNs), Long Short-Term Memory (LSTM) networks, Gated Recurrent Units (GRUs), and Transformer models, with training window sizes ranging from 3 weeks to 15 weeks.
The window size is a crucial hyperparameter for these models. Thus, we evaluated the models using window sizes of 3, 5, 7, 10, 12, and 15 weeks, and the results are presented in Fig. 5. The findings align with our expectations: a window that is too short does not adequately capture the temporal dependencies of the data, whereas an excessively long window introduces redundancy. A window size of seven weeks strikes an optimal balance and is well-suited for this forecasting task.
Figure 6: MASE comparison with Recurrent Neural Networks (RNNs), Long Short-Term Memory (LSTM) networks, Gated Recurrent Units (GRUs), and Transformers on the number of cases, deaths, and hospitalizations for the following 1 - 4 weeks.
The predictions for one-, two-, three- and four-week forecast are shown in Fig. 6. In comparison to PINNs, the forecasts generated by RNNs, LSTMs, GRUs, and Transformers exhibit noticeable fluctuations, indicating a higher degree of overfitting. This contrast underscores the advantage of incorporating physical information into PINNs, which contributes to more stable and reliable predictions. A quantitative comparison is provided in Tab. 3, where PINNs consistently outperform all other prominent neural network models commonly used for time series forecasting, both in terms of MASE for point forecasting and WIS for quantile forecasting. Despite architectural advances in RNNs, LSTMs, GRUs, and Transformers, none of these models demonstrates substantial performance gains on this task. These findings emphasize the importance of integrating physics-based domain knowledge into neural network models for improved generalization and predictive accuracy.
@@ 4 Discussion and Summary
Epidemic forecasting can traditionally be framed as an inverse-forward problem, wherein model parameters must first be inferred before generating future predictions. In many scenarios, the mechanisms underlying disease transmission are partially understood and represented using compartmental models in the form of ODEs. However, several model parameters remain unknown and must be estimated by fitting the model to observational data. Once the parameters are identified, these compartmental models can be employed to forecast future trends in disease spread and evaluate the impact of public health interventions or the emergence of new variants.
Despite their utility, traditional mathematical models suffer from inherent limitations due to model misspecification. For example, they often assume fixed transition rates, which makes it challenging to adapt to evolving conditions, such as changes in transmission dynamics caused by new variants. Moreover, forecasting is not feasible unless all relevant parameters are accurately estimated. In contrast, data-driven neural network models have emerged as powerful tools for addressing complex social and scientific problems in recent years. However, conventional neural networks typically require large volumes of training data and labeled examples to perform effectively, and they often disregard prior domain knowledge. As a result, machine learning models may produce unrealistic or unreliable forecasts, particularly in extrapolative scenarios, due to their lack of grounding in the underlying physical or epidemiological processes.
In this work, we employed PINNs to integrate epidemiological knowledge, as described by compartmental models, with observed data to enhance the accuracy of infectious disease forecasts. Our results demonstrate that the proposed approach offers several key advantages that contribute to its superior performance. First, the PINN framework leverages neural networks to model complex and non-linear patterns inherent in infectious disease dynamics, enabling the system to capture intricate relationships that traditional models may miss. Second, PINNs incorporate domain-specific knowledge by embedding the compartmental models into the loss function. This integration ensures that the learned representations conform to the underlying principles and constraints of disease transmission. Moreover, the inclusion of mechanistic models in the training process helps mitigate overfitting, a common issue when relying solely on data-driven machine learning approaches. By constraining the learning process with known epidemiological laws, PINNs achieve a balanced trade-off between flexibility and generalizability.
@@
Additionally, PINNs can be employed to estimate key epidemiological parameters—such as transmission and recovery rates – directly from data, thereby enhancing the model’s adaptability to evolving disease dynamics. In particular, PINNs treat these parameters as outputs of the neural network, allowing them to vary over time. This enables the model to capture temporal changes in transmission behavior, such as those driven by viral mutations or shifts in public health interventions. Furthermore, PINNs are capable of continuously incorporating newly acquired data, allowing for real-time model updates and progressively improved forecasts as more information becomes available during an ongoing outbreak. This adaptability makes PINNs particularly well-suited for real-world epidemic surveillance and response. In addition, PINNs can integrate heterogeneous data sources – including social, environmental, and demographic factors – offering a more comprehensive and holistic framework for infectious disease forecasting.
@@
Recently, several PINNs-based models have been developed for infectious disease forecasting, including applications to COVID-19 [66, 67, 68] and measles [69]. Compared to these existing approaches, our proposed model incorporates a more complex compartmental framework with nine state variables embedded into the neural network architecture. This significantly increases the challenge of ensuring consistency with the underlying physics described by the ODE system, thereby demonstrating the flexibility and robustness of our model. Notably, our model achieved better performance metrics – Normalized Root Mean Squared Error (NR1 and NR2) and Normal Deviation (ND), as defined in [67] – compared to the Epidemiologically-Informed Neural Networks (EINNs) [67] in forecasting COVID-19 deaths, which is the only predicted state variable in EINNs. Specifically, our model achieved 0.53 (EINNs: 0.54) on NR1, 0.15 (EINNs: 0.24) on NR2, and 0.33 (EINNs: 0.38) on ND, while simultaneously predicting two additional state variables. Furthermore, while prior models primarily offer point forecasts, we extend our approach to quantile forecasts, which allows us to quantify uncertainty in the predictions.
Despite the strengths of the vanilla PINN framework in handling noisy and limited data, it lacks a built-in mechanism for quantifying the uncertainty of its predictions in the presence of input noise. A well-established method for regression with uncertainty estimation is Gaussian Process Regression (GPR), which, rooted in the Bayesian framework, provides a principled way to quantify uncertainty in the solution of differential equations. However, GPR is typically limited to small, linear problems due to its high computational complexity, which scales poorly with large datasets. A promising direction to address this limitation involves the use of Bayesian PINNs, as proposed by [38]. These models combine the physics-informed learning capability of PINNs with the probabilistic nature of Bayesian inference, allowing for uncertainty-aware solutions even in ill-posed settings, such as those with incomplete or under-specified boundary conditions. Future research should investigate the applicability and effectiveness of Bayesian PINNs in the context of infectious disease forecasting, particularly in scenarios involving high uncertainty, evolving dynamics, or limited data availability.
@@
This work was supported by National Institute of Health grants R21HL168507 and NSF SCH Award Number:2406212. High-performance computing resources were provided by The Georgia Advanced Computing Resource Center (GACRC) at the University of Georgia.